kingston:

Dr. Fritz, a cat and dog veterinarian, examined 40% cats this week at his clinic. If Dr. Fritz examined 22 cats, how many cats and dogs did he examine altogether?

Gdeinward:

So here in this problem 40 percent is equivalent to 22 animals regardless of the distinction. So we can say that 80 percent is equivalent to 44 animals, If we double the percentag e we double the number. So the goal here is to find the total number of animals or 100 percent.

Gdeinward:

So now we know that 100 percent minus the 80 percent we found is 20 percent, or half of 40 percent.

Gdeinward:

Since we know that 40 percent is equal to 22 cats, we can subtract half the percentage and half of 22 to get 20 percent equals 11 animals. Now we know the total amount of animals is 55, or a hundred percent. So if theres 55 total animals and 22 cats, we subtract, 55 - 22 to get 33 dogs.

Gdeinward:

this can be put into equation form as well: 40% = 22 40% * 2 = 22 * 2 80% = 44 40% / 2 = 22 / 2 20% = 11 so 40% = 22 + (40%=22) + (20% = 11) = 100 percent = 55 55 - 22 cats = 33 dogs.